# Stability analysis of power-law cosmological models

### v1

@@ III Re-examining the stability of power-law solutions
$X^{(2)}=-\frac{\lambda}{\sqrt{6}}+D_{1}\,t^{\frac{7}{10}}+C_{1}\,t^{\frac{7}{10}}+\frac{20\,D_{1}^{2}\,\sqrt{30}}{119}\,t^{\frac{7}{5}}-\frac{7D_{1}\,D_{2}}{t^{\frac{3}{10}}}$
In the second-order terms, note the presence of $10\,D_{2}^{2}/t^{3}=(-D_{2}/p)^{2}(p/t^{3})$, which is the second term in the Taylor series expansion of $H_{p}(t-D_{2})$. Hence, we can argue that this term is not physical. Now, $\Delta^{(2)}H$ contains terms growing faster than $1/t$ which are $C_{2}/t^{2}$ and $-3D_{1}\,D_{2}\sqrt{30}/(85\,t\,t^{(3/10)})$. Here, $C_{2}/t^{2}$ corresponds to the first-order term in the Taylor expansion of $H$. Hence, we can conclude that if time translation alone explains the growing mode in $H$, then $H$ can be approximated to
@@
Like, as shown here, $H$ is closer to the solution $H_{c}$ compared to $H_{p}(t-D_{2})$, makes us guess whether the solution could be explained by a better time-translated solution. We don’t think that can happen because we don’t think the term $-3D_{1}\,D_{2}\sqrt{30}/(85\,t\,t^{(3/10)})$ will be eliminated by a higher order term. In first order there weren’t a growing mode other than $D_{2}/t^{2}$ which could be explained by time translation and is the reason why incorporating the terms that lead to time translation in second-order significantly made the solution approach the time translated solution. We don’t think this will happen at subsequent orders.
That is, the presence of $-3D_{1}\,D_{2}\sqrt{30}/(85\,t\,t^{(3/10)})$ makes the problem non-trivial. Similarly, $X^{(2)}$ also contains a growing mode that is physical. In the case of $X$, any term with $t$ raised to a negative power can be considered a growing term. However, the growing term diverges so slowly that it can be neglected because, much before the growing term becomes significant, the solution supposedly attains a big crunch. Our results clearly say that the contracting solutions are not attractors.
Now, we resort to numerical computation to see whether our conclusion is correct when the complete solutions are considered. In 3 and 3, we have compared the analytical expression $H^{(2)}/H_{c}$ and $X^{(2)}/X_{p}$ (the solid curve) with the numerical computation $H_{num}/H_{c}$, the dotted curves. Note that $X_{p}(t+\delta)=X_{p}(t)=X_{p}$. Here, we have taken only one curve, that is, with just one set of initial conditions. For numerical results, we had to resort to extremely high accuracy (in maple setting $absolute\,error=10^{(-23)}$ and $relative\,error=10^{-28}$). The numerical calculation breaks as shown with different coloured curves in the plot. We stitch the curves by setting a new initial condition that matches the analytical solution. Our argument here is that if the solution is an attractor, the curves would have moved towards the value, irrespective of the slight deviations it may encounter. Since the intermittent initial conditions come from the analytical solutions, the value of $D_{1}$, $D_{2}$, $C_{1}$, or $C_{2}$ won’t change. Hence, with this result, we argue that the Power-Law contraction is a repeller, as suggested by numerical results.
Figure 2: The evolution of $H$ for a single set of initial conditions. The curve is stitched together, as shown by curves coloured differently. The analytic solution is also plotted along. For an attractor solution, the curve must evolve to the value $1$. Instead, the curve deviates away. There isn’t a perfect match between analytic and numerical results, though both indicate the solution to be a repeller. (Note that $H\_appx$ is $H^{(2)}$. The numerical method is also mentioned in the legend)
Figure 3: The evolution of $X$ for a single set of initial conditions. The curve is stitched together, as shown by curves coloured differently. The analytic solution is also plotted along. For an attractor solution, the curve must evolve to the value $1$. And, that is what we are getting, indicating $X$ is an attractor in time. (Note that $X\_appx$ is $X^{(2)}$. The numerical method used is also mentioned in the legend)
The inflationary solution for $\lambda<6$ also shows instability in the second order; however, before the growing terms become significant, we get a sufficient amount of inflation as $p\gg 1$ for this scenario (for large values of $p$, the scale factor expands faster, and we get a sufficient amount of inflation before the second order growing modes become significant to affect the evolution). On the other hand, for large values of $\lambda$, that is, $p\ll 1$, the growing terms kick in before we have sufficient contraction (because if p is very small, the Universe evolves slowly). With much importance, we would like to point out the evolution of $N$ in time 28; here, one can see that the evolution of $N$ in $t$ is not always an attractor and must be taken into account while doing the stability analysis.
@@
Our result is of great importance in bouncing models. So, a power-law contraction can never be an attractor in this setup. This poses serious problems for the ekpyrotic scenario.
III.1 Attractor nature of X in N
@@
By rewriting the differential equation 10 and solving, we can obtain a relation between $N$ and $X$. This is a complete solution, and the integration constant, $C_{3}$, could be absorbed into N. If any complex value arises from the logarithmic terms, then $C_{3}$ can be split up into two parts, one to cancel any complex value that arises from logarithmic terms and the other to incorporate the initial deviation in X, which could be absorbed into N. This, together with the fact that N approaches $\infty$- ∞ as X approaches $X_{p}$ shows that $X$ is an attractor in $N$.
@@
$N=2\left(\ln\left(6X-2\sqrt{6}\sqrt{5}\right)\frac{1}{14}-\left(3\ln\left(X^{2}-1\right)-2\sqrt{6}\sqrt{5}\mbox{Tanh}^{-1}\left(X\right)\right)\frac{1}{64}\right)+C_{3}$ (31)
@@ IV Conclusion
One of the primary reasons we think the Big Bang cosmology is incomplete is that the model suffers from cosmological puzzles, among which the fine-tuning problems are most important. For this very reason, we want the early Universe models, whether Inflation or bounce, to be attractors. Power-law inflation is one of the most important inflation models because of its simplicity and ability to explain observations, as well as because this model seeds many different cosmological models. There are plenty of cosmological models in the literature motivated by power-law inflation [5, 21, 37, 22, 38]. Hence, the stability analysis of this model is of utmost importance. And it is a topic that has been extensively studied. The study mainly focuses on converting the differential equations in time to differential equations in $N$, defining new quantities $X=\dot{\phi}/(\sqrt{6}H)$ and $Y=\sqrt{V}/(\sqrt{3}H)$, and obtaining the conditions for the stability of these quantities. In this process, we failed to check whether the stability of $X$ and $Y$ ensures the stability of the relevant background quantities. In this paper, we study the cosmological evolution in time and for the relevant background quantities, this was explored previously, see [29, 2] for details. We found out that the conditions turn out to be the same for inflation (expanding Universe) even though the form of evolution differs. However, they drastically differ for a contracting Universe, which means the bouncing models’ stability must be reanalyzed. We conclude that power-law contraction is always a repeller, irrespective of the value of the parameter $\lambda$. Also, we note that the argument of time shift making the growing mode nonphysical [29] at first order won’t save power-law contraction from being unstable; at higher orders, there are growing modes that are physical. We explicitly show the case in second order. Hence, ekpyrotic bounces and other models built based on power-law models must be taken cautiously. Most importantly, our paper shows that how stability analysis is done has to be re-examined. This makes our paper important in similar studies. We want to add to our statement that our analysis doesn’t point out that the previous calculations were wrong; instead, it strengthens the previous calculations and results. As pointed out in earlier works $X$ and $Y$ are attractors in time for $\lambda^{2}>6$. But, by pointing out that the evolution of $H$ and $\phi$ in time are the ones we must look into and they diverge, our conclusions differ from previous conclusions.
Acknowledgements and Authors contribution

### v2

@@ III Stability of power-law solutions
$X^{(2)}$ $=\sqrt{\frac{7}{6}}-D_{1}(-t)^{1/7}$
$+\left(\frac{7D_{1}^{2}(-t)^{2/7}\sqrt{6}\sqrt{7}}{4}-C_{1}(-t)^{1/7}-\frac{D_{1}D_{2}}{2(-t)^{6/7}}\right)$
Since, while solving for the integration constants, we consider the first-order solution and second-order solution separately, $D_{1}$ and $C_{1}$ are two different constants that can be determined from initial conditions.
We would like to stress the fact that the $FRW$ equations in the conventional form require two integration constants to define the system completely. The equations we use also require two integration constants. Hence, we can confidently say that no additional constraints are there in the system.
In the second-order terms in the Hubble parameter, note the presence of $7\,D_{2}^{2}/2t^{3}=(-D_{2}/p)^{2}(p/t^{3})$, which is the second term in the Taylor series expansion of $H_{p}(t-D_{2})$. Hence, we can argue that this term is not physical. Now, $\Delta^{(2)}H$ contains terms growing faster than $1/t$ which are $C_{2}/t^{2}$ and ${3D_{1}D_{2}\sqrt{42}}/{(14(-t)^{13/7})}$. Here, $C_{2}/t^{2}$ is equivalent to the first-order term in the Taylor expansion of $H$. Hence, we can conclude that if time translation alone explains the growing mode in $H$, then $H$ can be approximated to
@@
As shown here, $H$ is closer to the solution $H_{c}$ compared to $H_{p}(t-D_{2})$, making us guess whether the solution could be explained by a better time-translated solution. We don’t think that can happen because it is highly improbable that the term ${3D_{1}D_{2}\sqrt{42}}/{(14(-t)^{13/7})}$ will be eliminated by a higher-order term, unlike this term here in the second order, among first-order terms there wasn’t a growing mode other than $D_{2}/t^{2}$, $D_{2}/t^{2}$ could be explained by time translation, and this is the reason why incorporating $C_{2}$ appearing in second-order perturbed terms, along with $D_{2}$ explains the solution better. Hence, this will not happen with subsequent orders. That is, the presence of ${3D_{1}D_{2}\sqrt{42}}/{(14(-t)^{13/7})}$ in $H$ makes the problem non-trivial. Similarly, $X^{(2)}$ also contains a growing mode that is physical. In the case of $X$, any term containing time ($t$ ) raised to a negative power can be considered a growing term. $X$ is a better variable to capture the stability of the solution because, in the case of $H$ or, for that matter, any other variable with the particular solution that is a function of time, is ailed by the uncertainty in figuring out the correct time translation. New higher-order perturbed terms may be added to the time translation when exploring higher order terms.
We also show numerically that the system is unstable when evolving in $t$. See 1, where $X/X_{p}$ obtained both numerically (We solve equations  21b numerically.) and analytically are plotted against time for different initial conditions, see 1 for the values of the initial conditions we choose and the values of respective integration constants. The plot clearly demonstrates that the system is unstable when it evolves in time. If it was an attractor, the curves would have evolved towards $1$. But, here, we can see that the curves are repelled away from $1$. In the next section, we show that $X$ is an attractor in $N$. We discuss the implications of these confusing and exciting results in the conclusion. Our results clearly say that the contracting solutions are not attractors.
For numerical results, we had to resort to extremely high accuracy (in maple, setting $\mbox{absolute}\,\mbox{error}=10^{(-22)}$, $\mbox{relative}\,\mbox{error}=10^{-28}$ and Digits, the precision in number of decimal digits, to be $300$). The numerical calculation breaks as the step size effectively reaches zero [^4] We tried to solve the differential equations numerically using $RK4$, $RK6$ and $RKf45$ adaptive step size method in $C++$, Here, we used arbitrary precision libraries such as BOOST multi-precision lybrary , mpfr lybrary. et cetra. However, we figured out that nothing can match the advanced algorithms of maple, both in speed and precision. Also, whether to use $C++$ or maple depends on the specific problem. It was found that for the evolution in $N$, $C++$ offers a better choice which also helps in understanding the problem better. Here, we would like to point out that since the numeric computation breaks by reaching step size effectively zero even with such high precision, we must not resort to numeric computation. It can only be used to guide us in our study. For this reason, we think it’s the analytic result that correctly depicts the scenario. Also, note that the integral of $H$ with respect to $t$ gives $N$ and is also a repeller in time. We hope this explains our confusing result. On the other hand, the inflationary solution with $\lambda^{2}<6$ is found to be stable. The second-order perturbative analysis also shows no deviation from the standard picture.
@@
So, a power-law contraction can never be an attractor in this setup. This poses serious problems for the ekpyrotic scenario.
Figure 1: The evolution of $X$ for different initial conditions. We plot $X(t)/X_{p}$ obtained numerically ($X_{num}/X_{p}$) and analytically ($X^{(2)}/X_{p}$) $vs$ Time in $\mbox{Time in}\;1/M_{p}$. See 1 for initial conditions and other details. We used the $RKf45$ adaptive step size method available with maple for numerical calculations. Due to the inability to attain sufficient accuracy, the numerical curves break intermittently
Curve $\Delta H(t_{i})$ $\Delta X(t_{i})$ $D_{1}$ $D_{2}$ $C_{1}$ $C_{2}$
1 $-2.86\times 10^{-34}$ $5.40\times 10^{-2}$ $-2.80\times 10^{-6}$ $-2.53\times 10^{28}$ $1.68\times 10^{-6}$ $3.36\times 10^{27}$
2 $2.86\times 10^{-32}$ $-1.08\times 10^{-1}$ $5.59\times 10^{-6}$ $7.86\times 10^{28}$ $6.63\times 10^{-6}$ $2.12\times 10^{28}$
3 $-2.86\times 10^{-32}$ $-5.40\times 10^{-2}$ $2.80\times 10^{-6}$ $-3.57\times 10^{27}$ $1.72\times 10^{-6}$ $3.69\times 10^{27}$
4 $2.86\times 10^{-32}$ $1.08\times 10^{-1}$ $-5.59\times 10^{-6}$ $-2.14\times 10^{28}$ $6.79\times 10^{-6}$ $1.12\times 10^{28}$
5 $-2.86\times 10^{-35}$ $2.16\times 10^{-1}$ $-1.11\times 10^{-5}$ $-1.00\times 10^{29}$ $2.69\times 10^{-5}$ $5.33\times 10^{28}$
6 $2.86\times 10^{-36}$ $-2.16\times 10^{-1}$ $1.12\times 10^{-5}$ $1.00\times 10^{29}$ $2.69\times 10^{-5}$ $5.33\times 10^{28}$
Table 1: Table showing the initial conditions $\Delta H(t_{i})=H(t_{i})-H_{p}(t_{i})$ and $\Delta X(t_{i})=X(t_{i})-X_{p}(t_{i})$ chosen at random but maintaining $H(t_{i})$ and $X(t_{i})$ near to $H_{p}(t_{i})$ and $X_{p}$, and the corresponding values of integration constants for each curve in 1.
III.1 Stability of $X$ and $H$ in number of e-foldings
@@
By rewriting the differential equation 11 and solving, we can obtain a relation between $N$ and $X$. This is a complete solution, and the integration constant, $C_{3}$, could be absorbed into $N$. If any complex value arises from the logarithmic terms, then $C_{3}$ can be split up into two parts, one to cancel any complex value that arises from logarithmic terms and the other to incorporate the initial deviation in $X$, which could be absorbed into $N$. This, together with the fact that $N$ approaches $\infty$- ∞ as $X$ approaches $X_{p}$, shows that $X$ is an attractor in $N$.
$N+C_{3}=-\frac{\ln\left(X+1\right)}{\sqrt{6}\,\lambda-6}+\frac{\ln\left(X-1\right)}{\sqrt{6}\,\lambda+6}+\frac{12\ln\left(\sqrt{6}\,\lambda+6X\right)}{\left(\sqrt{6}\,\lambda+6\right)\left(\sqrt{6}\,\lambda-6\right)}$ (29)
Now, we know that the slow-roll parameter $\epsilon_{1}$, given by $-\dot{H}/H^{2}$ can be written in terms of $X$ alone. That is, $\epsilon_{1}=3X^{2}$. Now, as $X$ approaches $X_{p}$, $\epsilon_{1}$ approaches $\lambda^{2}/2$. It’s not surprising that this happens when $N\rightarrow-\infty$. But at this point, we have $\epsilon_{1}=\infty/\infty$ as both $\dot{H}$ and $H$ goes to $\infty$.
Let us also do an order-by-order analysis in $N$ for $X$ and $H$. Here we perturb the equations 10 and 11. We get the differential equations for $X$ to be
$0=$ $\left(\frac{d}{dN}\Delta X-\frac{\lambda^{2}\Delta X}{2}+3\Delta X\right)\epsilon+\left(\sqrt{6}\,\lambda\,\Delta X^{2}+\frac{d}{dN}\Delta^{(2)}X-\frac{\lambda^{2}\Delta^{(2)}X}{2}+3\Delta^{(2)}X\right)\epsilon^{2}$
@@
$+$ $\left(2\sqrt{6}\,\lambda\,\Delta X\,\Delta^{(2)}X+\frac{d}{dN}\Delta^{(3)}X-3\Delta X^{3}-\frac{\lambda^{2}\Delta^{(3)}X}{2}+3\Delta^{(3)}X\right)\epsilon^{3}\hfill$ (30)
and the equations for $H$ to be
$0=\left(\frac{d}{dN}\Delta H+\frac{\lambda^{2}\Delta H}{2}-\frac{2\sqrt{6}\,\Delta X\,e^{-\frac{N\lambda^{2}}{2}}}{\lambda}\right)\epsilon$
$+\left(\frac{d}{dN}\Delta^{(2)}H+\frac{6\Delta X^{2}\,e^{-\frac{N\lambda^{2}}{2}}}{\lambda^{2}}-\sqrt{6}\,\lambda\,\Delta X\,\Delta H+\frac{\lambda^{2}\Delta^{(2)}H}{2}-\frac{2\sqrt{6}\,\Delta^{(2)}X\,e^{-\frac{N\lambda^{2}}{2}}}{\lambda}\right)\epsilon^{2}$
$+\left(\frac{d}{dN}\Delta^{(3)}H-\sqrt{6}\,\lambda\,\Delta X\,\Delta^{(2)}H-\sqrt{6}\,\lambda\,\Delta^{(2)}X\,\Delta H+\frac{\lambda^{2}\Delta^{(3)}H}{2}+3\Delta X^{2}\,\Delta H-\frac{2\sqrt{6}\,\Delta^{(3)}X\,e^{-\frac{N\lambda^{2}}{2}}}{\lambda}+\frac{12\Delta X\,\Delta^{(2)}X\,e^{-\frac{N\lambda^{2}}{2}}}{\lambda^{2}}\right)\epsilon^{3}$ (31)
Now we have the solution for $X$ till the third order given by
$\Delta X$ $=c_{1}e^{\frac{(\lambda^{2}-6)N}{2}},$ (32)
$\Delta^{(2)}X$ $=-\frac{2c_{1}^{2}\sqrt{6}\,e^{(\lambda^{2}-6)N}\lambda}{\lambda^{2}-6}+e^{\frac{(\lambda^{2}-6)N}{2}}c_{2},$ (33)
$\Delta^{(3)}X$ $=e^{\frac{(\lambda^{2}-6)N}{2}}c_{3}\quad-\quad\frac{c_{1}\left(e^{\frac{(\lambda^{2}-6)N}{2}}\right)^{2}}{\left(\lambda^{2}-6\right)^{2}}\;\;\bigg{(}4c_{2}\lambda^{3}\sqrt{6}-27\,e^{\frac{(\lambda^{2}-6)N}{2}}c_{1}^{2}\lambda^{2}-24\lambda c_{2}\sqrt{6}+18\,e^{\frac{(\lambda^{2}-6)N}{2}}c_{1}^{2}\bigg{)},$ (34)
and $H$, till third order given by
$\Delta H$ $=\left(\frac{4\sqrt{6}\,c_{1}e^{\frac{(\lambda^{2}-6)N}{2}}}{(\lambda^{2}-6)\lambda}+c_{4}\right)e^{-\frac{N\lambda^{2}}{2}},$ (35)
$\Delta^{(2)}H$ $=e^{-\frac{N\lambda^{2}}{2}}c_{5}-\frac{2\,e^{\frac{(\lambda^{2}-6)N}{2}}e^{-\frac{N\lambda^{2}}{2}}}{\lambda^{2}(\lambda^{2}-6)}\quad\left(-c_{4}\lambda^{3}c_{1}\sqrt{6}+3\,e^{\frac{(\lambda^{2}-6)N}{2}}c_{1}^{2}-2\lambda c_{2}\sqrt{6}\right),$ (36)
$\Delta^{(3)}H=\frac{{2\,{\mathrm{e}}^{-\frac{N\,\lambda^{2}}{2}}}}{2\left(\lambda^{2}-6\right)^{2}{\lambda^{2}}}\left(2\left(\left(c_{1}c_{5}+c_{2}c_{4}\right)\lambda^{2}+2c_{3}\right)\lambda\left(\lambda^{2}-6\right)\sqrt{6}\,{\mathrm{e}}^{\frac{\left(\lambda^{2}-6\right)N}{2}}+8\sqrt{6}\,c_{1}^{3}{\mathrm{e}}^{\frac{3\left(\lambda^{2}-6\right)N}{2}}\lambda\right.$
$\left.-3\left(\lambda^{2}-6\right)\left(c_{1}\left(c_{1}c_{4}\lambda^{2}+4c_{2}\right){\mathrm{e}}^{\left(\lambda^{2}-6\right)N}-\frac{\lambda^{2}c_{6}\left(\lambda^{2}-6\right)}{3}\right)\right).$ (37)
The above calculations clearly show that $H$ and $X$, and hence the other background quantities are attractors in $N$. Note that the terms that appear to grow in the expression for $H$ faster than $H_{p}(N)={2\exp\left(-{N\lambda^{2}}/{2}\right)}/{\lambda^{2}}$ can be explained by a translation in $N$. These terms are similar to the time translation terms we discussed earlier. Now, in time, the background quantities are not attractors. This is confusing, and we think this happens because, in time, the Universe doesn’t follow a contracting path indefinitely, and $N$ doesn’t take larger and larger negative values following the power law trajectory. That is, $-N$ might not be a monotonically increasing function in time. Instead, there might be an alternate path in which the Universe evolve. We can see from the expression for $H$ that $N$ is a repeller in time. However, we must admit that we haven’t fully understood what is happening here.
@@ IV Conclusion
One of the primary reasons the Big Bang cosmology is incomplete is that the model suffers from cosmological puzzles, among which the fine-tuning problems are most important. For this very reason, we want the early Universe models, whether Inflation or bounce, to be attractors. Power-law inflation is one of the most important inflation models because of its simplicity and ability to explain observations, as well as because this model seeds many different cosmological models. There are plenty of cosmological models in the literature motivated by power-law inflation [6, 22, 42, 23, 43]. Hence, the stability analysis of this model is of utmost importance. And it is a topic that has been extensively studied. The study mainly focuses on converting the differential equations in time to differential equations in $N$, defining new quantities $X=\kappa\dot{\phi}/(\sqrt{6}H)$ and $Y=\kappa\sqrt{V}/(\sqrt{3}H)$, and obtaining the conditions for the stability of these quantities. In this process, one must also check that the stability of $X$ and $Y$ and other background quantities in $N$ ensures the stability of the relevant background quantities in time.
In this paper, we study the cosmological evolution in time for the relevant background quantities up to the second order; this was explored previously to the first order; see [2, 3] for details. We found out that the conditions for a stable evolution turn out to be as expected for inflation. However, they drastically differ for a contracting Universe, which means the bouncing models’ stability must be reanalyzed. In the second order, we explicitly showed that there are physical growing modes in $X$ and $H$, which means the system is unstable in time. We conclude that power-law contraction is always a repeller, irrespective of the value of the parameter $\lambda$. Also, we note that the argument of time shift making the growing mode nonphysical [2] at first order won’t save power-law contraction from being unstable; at higher orders, there are growing modes that are physical. Hence, ekpyrotic bounces and other models built based on power-law models must be taken cautiously. Interestingly, we got the background quantities to be attractors in the number of e-foldings and repellers in time. This is confusing, and we must admit that we don’t fully understand what is happening here. Because if a solution, when looked at from one variable, is an attractor, it must be an attractor when looked at from any other variable that is monotonically increasing with respect to the first variable. However, it must be noted that in the case of contracting evolution, the number of e-foldings diverges away from the expected evolution and is a repeller in time. This could be one of the possible reasons why we are getting the background quantities to be repellers in time when, at the same time, they are attractors in $N$. We think $N$ is not evolving to sufficiently large negative values that the background quantities are attractors. Also, it could be possible that, in time, the solution is a repeller for a short period, and it will evolve back to the power-law solution. However, looking at the form of the second-order terms, it is highly unlikely. This result, if true, has significant consequences. Because it hints that all temporal variables cannot be considered equivalent. Most importantly, our paper shows that the way in which stability analysis is done has to be re-examined. We want to add to our statement that our analysis doesn’t point out that the previous calculations were wrong; instead, it strengthens the previous calculations and results. As pointed out in earlier works, we got background quantities are attractors in the number of e-foldings during contraction for $\lambda^{2}>6$. But, by pointing out that the evolution of $H$ and $\phi$ in time are the ones we must look into, and they diverge, our conclusions differ from previous conclusions.
Acknowledgements and Author Contributions